Feedback 9SeeHint The ratio of hydrogen atoms (H) to carbon atoms (C) in the Suns atmosphere is approximately 2,400 to 1. It is reasonable to assume this also applies to molecular clouds. Radio observations show that a molecular cloud contains 100 Mo of carbon monoxide (CO) in a giant molecular cloud How much carbon is contained in the cloud? Carbon represents 3/7 of the mass of a CO molecule. Mo 943 How much hydrogen is contained in the cloud? Remember that hydrogen weighs about 1/12 of carbon 8572 Mo
Answer #1

Atoms ratio H:C = 2400:1

Mass ratio = 2400*1:1*12 = 200:1

Carbon contained in cloud = 43

Hence H contained in cloud = 43*200

= 8600M